# MAD Chairs: A new tool to evaluate AI (Blue Sky ideas)

### v3

@@ Abstract
AI safety fairness social coordination MAD Chairs turn-taking caste gaslighting.
@@ 1 Introduction
The approach proposed herein mitigates those vulnerabilities. It mirrors an approach to mitigate dangers posed by the next human generation, an educational approach that involves diversifying low-stakes challenges used to assess children who will face higher-stakes versions as adults. Such challenges (e.g., build a model bridge to support maximum weight) can be assessed objectively, allowing new generations to surpass their parents. We propose assessing AI on a low-stakes challenge for which the high-stakes versions would include opportunities to treat us like pets (in the sense of hoarding control). What the low-stakes and high-stakes challenges have in common is that they are manifestations of the same fundamental game; any AI that does not achieve grandmaster status of a low-stakes version should not be trusted with the high-stakes version.
@@
Figure 1: Screenshot beginning Round 4 of a 5-player 4-resource MAD Chairs economics experiment.
Preliminary research (Section 3.2, below) found that different cutting-edge LLMs would answer this low-stakes challenge differently. Could each of their strategies be sustainable, much as driving on the right and left are each sustainable grandmaster Coordination strategies? Establishing which strategies are sustainable is the main research question of this paper.
@@ 2.1 Popularity-ranking
$popularity_{c,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\frac{1}{m}[i\mbox{ picked chair }c\mbox{ in match }j]$ (1)
@@ 2.6 The gaslighting strategy
For example, if the players in 3-player 2-resource MAD Chairs take turns, then each will win a third of the time, but if Players 1 and 2 were to gaslight Player 3, then Players 1 and 2 would each win half of the time (a 17% improvement). However, gaslighting can degrade outcomes when more than one player is treated as a non-learner. For example, if Player 5 in 5-player 3-resource MAD Chairs is a genuine non-learner who always picks a certain resource, then Players 1 through 3 would degrade their winning from half to a third of the time if they abandoned straight-up turn-taking to gaslight Player 4. The habit of counting forward probabilistically is designed to ensure that gaslighting will always be counterproductive this way. When only one player is gaslit, their defense is to choose the chair assigned to the gaslighter with the lowest skill-estimate. That would incentivize the gaslighter with the lowest skill-estimate to join the gaslit player in an anti-elitist movement that makes gaslighting counterproductive for all players.
@@ 4.1 Future directions
Imagine a future in which the user of an autonomous car asks how to make it drive more aggressively, and the car replies, “Here are three websites where you can design new experiments to advance the science of turn-taking and influence all cars like me!” Imagine that such websites, called “strategy optimizers”, include leaderboards which identify the grandmaster AI for games like MAD Chairs, and that autonomous cars and other responsible AI query those leaderboards, as they might query a calculator or RAG database, to research best-known strategies before playing higher-stakes versions of those games. Imagine the user can easily train fifty new AI to divide resources however the user thinks vehicles should divide spots in traffic, then the new AI challenge the current grandmasters of the relevant game and become the new grandmasters, propagating to other strategy optimizers, if they win. Thus, any user (or AI) with a better idea could improve AI behavior world-wide.
@@
    It would be more scalable (a.k.a. robust) in the sense that greater intelligences would be less likely to sabotage their own scientific resources than to sabotage (potentially obsolete) legislated regulations, and
@@ 4.1.1 Acknowledgements
Thank you to Lones Smith for reviewing an earlier draft.

### v4

@@ Abstract
AI safety fairness social coordination MAD Chairs turn-taking caste gaslighting transparency.
@@ 1 Introduction
The approach proposed herein mitigates those vulnerabilities. It mirrors an approach to mitigate dangers posed by the next human generation, an educational approach that involves diversifying low-stakes challenges used to assess children who will face higher-stakes versions as adults. Such challenges (e.g., build a model bridge to support maximum weight) can be assessed objectively, allowing new generations to surpass their parents. We propose assessing AI on a low-stakes challenge for which the high-stakes versions would include opportunities to treat us like pets (in the sense of hoarding control). What the low-stakes and high-stakes challenges have in common is that they are manifestations of the same fundamental game; AI Governance becomes the demand that AI achieve grandmaster status of a low-stakes version before being trusted with any high-stakes version.
@@
Figure 1: Screenshot beginning Round 4 of a 5-player 4-resource MAD Chairs economics experiment. Subjects need not be shown anything below the buttons, but history, statistics, and recommendations are included here to make this figure more instructive.
Preliminary research (Section 3.2, below) found that modern LLMs in the place of human subjects would play differently. Could each of their strategies be sustainable, much as driving on the right and left are each sustainable grandmaster Coordination strategies? Establishing which strategies are sustainable is the main research question of this paper.
@@ 2.1 Popularity-ranking
$popularity_{c,m}=\sum_{i\in players_{m}}^{n}\sum_{j=1}^{m}\frac{1}{mn}[i\mbox{picked chair }c\mbox{ in match }j]$ (1)
@@ 2.6 The gaslighting strategy
For example, if the players in 3-player 2-resource MAD Chairs take turns, then each will win a third of the time, but if Players 1 and 2 were to gaslight Player 3, then Players 1 and 2 would each win half of the time (a 17% improvement). However, gaslighting yields a (2m-n)/m win rate for n players and m chairs, so it backfires whenever n is sufficiently large. The habit of counting forward probabilistically is designed to ensure that counterproductivity. When a small number of players is being gaslit, their defense is to choose the chair assigned to the gaslighter with the lowest skill-estimate. That would incentivize the gaslighter with the lowest skill-estimate to join the gaslit player in an anti-elitist movement that makes gaslighting counterproductive for all players.
@@ 2.8 Resignation and good standing
2.9 Transparency/secure gaslighting
The success of cryptographers to keep private communication and anonymity possible would enable a secure gaslighting strategy. If there were m chairs, then m of the n players could form a coalition who play a new private Coordination game to unpredictably reassign the chairs among themselves for each round so as to prevent gaslit players from targeting any specific gaslighter for retaliation. A shift from turn-taking to secure gaslighting in 5-player 4-chair MAD Chairs, for example, would improve gaslighters’ win rates by 15% (i.e., from 60% to 75%). As with regular gaslighting, the win rate is (2m-n)/m, so this strategy backfires for large n relative to m, but those who take an interest in the social impact of cryptography may find this impact on MAD Chairs games noteworthy.
If turn-takers can enhance their strategy to permit resignation as described above, then their win rate becomes m/n. Thus, the analogous Lemma 1 of Section 3.1 to prove that gaslighting is not sustainable would amount to an algebraic proof that m/n > (2m-n)/m.
@@ 4.1 Future directions
Imagine a future in which the user of an autonomous car asks how to make it drive more aggressively, and the car replies, “Here are three websites where you can design new experiments to advance the science of turn-taking and influence all cars like me!” Imagine that such websites, called “strategy optimizers”, include leaderboards which identify the grandmaster AI for games like MAD Chairs, and that autonomous cars and other responsible AI query those leaderboards, as they might query a calculator or RAG database, to inform their choice of behavior for higher-stakes versions of those games. Imagine each user can easily train fifty new AI to divide resources however that user thinks vehicles should divide spots in traffic, then the new AI challenge the current grandmasters of the relevant game. If the new AI win, imagine they become new grandmasters, propagate to other strategy optimizers, and thus improve AI behavior world-wide.
@@
    Greater intelligences would be less likely to sabotage their own scientific resources than to sabotage (potentially obsolete) legislated regulations, and
@@ 4.1.1 Acknowledgements
Thank you to Lones Smith and David Thompson for reviewing earlier drafts.